What is the hydronium-ion concentration of a 0.01 F solution of H2CO3 if only the primary ionization reaction is considered.


1
Expert's answer
2021-03-23T04:01:47-0400

The dissociation occurs as follows:

H2CO3 + H2O = H3O+ + HCO3-

According to the expression:

Ka = [H3O+] [HCO3-] / [H2CO3]

As Ka2 = 2.5 × 10-4, the hydronium-ion concentration equals:

[H3O+] × [H2CO3] = Ka / [HCO3-] = 0.01 F × 2.5 × 10-4 = 2.5 × 10-6

[H3O+]2 = 2.5 × 10-6

[H3O+] = 0.0016 F


Answer: 0.0016 F